You need a 95% alcohol solution. On hand, you have a 60 mL of a 25% alcohol mixture. How much pure alcohol will you need to add to obtain the desired solution?

Answer by mananth(16946)

--- percent ---------------- Amount
Alcohol I 25 ---------------- 60 ml
Pure alcohol 100 ---------------- x ml
Mixture 95 ---------------- 60 + x ml

25 * 60 + 100 x = 95 ( 60 + x )
1500 + 100 x = 5700 + 95 x
100 x -95 x = -1500 + 5700
5 x = 4200
/ 5
x = 840 ml Pure alcohol

m.ananth@hotmail.ca